Consider the linear eigenproblem Ax=Ξ»x for the matrix



D=[1 1 2



2 1 1



1 1 3 ]



1. Solve for the largest (in magnitude) eigenvalue of the matrix and the corresponding eigenvector by the power method with π‘₯(0)T=[1 0 0]



2. Solve for the smallest eigenvalue of the matrix and the corresponding eigenvector by the inverse power method using the matrix inverse. Use Gauss-Jordan elimination to find the matrix inverse.

1
Expert's answer
2021-11-12T06:48:32-0500

1st Iteration;


"Dx_0=" "\begin{pmatrix}
 1&1&2 \\
 2&1& 1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 1 \\
 0\\
0
\end{pmatrix}=\begin{pmatrix}
 1 \\
 2 \\
1
\end{pmatrix}"



"x_1=\frac{1}{2}\begin{pmatrix}
 1 \\
 2\\
1
\end{pmatrix}=\begin{pmatrix}
 0.5\\
 1\\
0.5
\end{pmatrix}"



2nd Iteration



"Dx_1=\begin{pmatrix}
 1&1&2 \\
 2&1& 1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 0.5 \\
 1 \\
0.5
\end{pmatrix}=\begin{pmatrix}
 2.5 \\
 2.5 \\
3
\end{pmatrix}"


"x_2=\frac{1}{3}\begin{pmatrix}
 2.5 \\
 2.5 \\
3
\end{pmatrix}=\begin{pmatrix}
 0.83 \\
 0.83\\
1
 
\end{pmatrix}"



3rd Iteration



"Dx_2=\begin{pmatrix}
 1&1&2 \\
 2&1 & 1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 0.83 \\
 0.83 \\
1
\end{pmatrix}=\begin{pmatrix}
 3.67 \\
 3.5 \\
4.67
\end{pmatrix}"



"x_3=\frac{1}{4.67}\begin{pmatrix}
 3.67 \\
 3.5\\
4.67
\end{pmatrix}=\begin{pmatrix}
 0.79 \\
 0.75 \\
1
\end{pmatrix}"




4th Iteration



"Dx_3=\begin{pmatrix}
 1&1 & 2 \\
 2&1 & 1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 0.79 \\
 0.75\\
1
\end{pmatrix}=\begin{pmatrix}
 3.54 \\
 3.32 \\
4.54
\end{pmatrix}"




"x_4=\frac{1}{4.54}\begin{pmatrix}
 3.54 \\
 3.32 \\
4.54
\end{pmatrix}=\begin{pmatrix}
 0.78 \\
 0.73\\
1
 
\end{pmatrix}"




5th Iteration



"Dx_4=\begin{pmatrix}
 1&1 & 2 \\
 2&1 & 1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 0.78\\
 0.73 \\
1
\end{pmatrix}=\begin{pmatrix}
 3.51 \\
 3.29\\
4.51
\end{pmatrix}"




"x_5=\frac{1}{4.51}\begin{pmatrix}
 3.51 \\
 3.29 \\
4.51
\end{pmatrix}=\begin{pmatrix}
 0.78 \\
 0.73 \\
1
\end{pmatrix}"




6th Iteration



"Dx_5=\begin{pmatrix}
 1&1 & 2\\
 2&1& 1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 0.78 \\
 0.73 \\
1
\end{pmatrix}=\begin{pmatrix}
 3.51 \\
 3.29\\
4.51
\end{pmatrix}"




"x_6=\frac{1}{4.51}\begin{pmatrix}
 3.51 \\
 3.29 \\
4.51
\end{pmatrix}=\begin{pmatrix}
 

0.78 \\
 
 0.73\\
1
\end{pmatrix}"




"\therefore" The dominant eigenvalue "\lambda=4.51"

and the dominant eigenvector is


"\begin{pmatrix}
 0.78 \\
 0.73 \\
1
\end{pmatrix}"



D-1 by Gaus - Jordan Elimination


Augmenting "D" with a "3\u00d73" identity matrix


"\begin{bmatrix}
 1&1&2| 1&0&0 \\
 2&1&1|0&1&0 \\
1&1&3|0&0&1
\end{bmatrix}"


"\frac{1}{2}R_2-R_1\to\>R_2"


"\begin{bmatrix}
 1&1 & 2| 1&0&0\\
 0&\frac{-1}{2}& \frac{-3}{2}\>\>\>|-1&\frac{1}{2}&0\\
-1&1&3|0&0&1
\end{bmatrix}"



"R_3-R_1\to\>R_3"


"\begin{bmatrix}
 1&1&2\>\>\>| 2&0&0\\
 0&\frac{-1}{2} &\>\> \frac{-3}{2}\>\>\>\>\>|-1&\frac{1}{2}&0\\
0&0&\>\>1\>\>\>\>\>\>\>\>|-1&0&1
\end{bmatrix}"



"-2R_2-3R_3\to\>R_2"


"\begin{bmatrix}
 1&1&2 |1&0&0 \\
 0&1&0 |5&-1&-3\\
0&0&\>1\>\>\>\>\>|-1&0&1
\end{bmatrix}"



"R_1-2R_3\to\>R_1"


"\begin{bmatrix}
 1&1 & 0|3&0&-2\\
 0&1&0| 5&-1&-3\\
0&0&\>\>-1|-1&0&1
\end{bmatrix}"



"R_1-R_2\to\>R_1"


"\begin{bmatrix}
 1&0&\>\>\>\>0 | \quad-2&1&1\\
 0&1&0|\quad 5&-1&-3\\
0&0&1|-1&0&1
\end{bmatrix}"



D-1 is the right side of Augmented matrix


"D^{-1}=\begin{pmatrix}
 -2&1&1 \\
 5&-1&-3 \\
-1&0&1
\end{pmatrix}"




Solving for the largest eigenvalue of D-1


1st Iteration


"D^{-1}x_0=\begin{pmatrix}
 -2&1 & 1\\
 5&-1 & -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 1 \\
 0 \\
0
\end{pmatrix}=\begin{pmatrix}
 -2 \\
 5 \\
-1
\end{pmatrix}"




"x_1=\frac{1}{5}\begin{pmatrix}
 -2 \\
 5 \\
-1
\end{pmatrix}=\begin{pmatrix}
 -0.4 \\
 1\\
-0.2
\end{pmatrix}"



2nd Iteration


"D^{-1}x_1=\begin{pmatrix}
 -2&1 & 1\\
 5&-1 & -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 -0.4 \\
 1\\
-0.2
\end{pmatrix}=\begin{pmatrix}
 1.6 \\
 -2.4 \\
0.2
\end{pmatrix}"



"x_2=\frac{1}{2.4}\begin{pmatrix}
 1.6 \\
 -2.4 \\
0.2
\end{pmatrix}=\begin{pmatrix}
 0.67 \\
 -1 \\
0.08
\end{pmatrix}"



3rd Iteration


"D^-x_2=\begin{pmatrix}
 -2&1&1 \\
 5&-1 & -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 0.67 \\
 -1 \\
0.08
\end{pmatrix}=\begin{pmatrix}
 -2.25 \\
 4.08\\
-0.58
\end{pmatrix}"



"x_3=\frac{1}{4.08}\begin{pmatrix}
 -2.25 \\
 4.08 \\
-0.58
\end{pmatrix}=\begin{pmatrix}
 -0.55 \\
 1 \\
-0.14
\end{pmatrix}"




4th Iteration


"D^-x_3=\begin{pmatrix}
 -2&1 & 1 \\
 5&-1& -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 -0.55 \\
 1 \\
-0.14
\end{pmatrix}=\begin{pmatrix}
 1.96 \\
 -3.33 \\
0.41
\end{pmatrix}"




"x_4=\frac{1}{3.33}\begin{pmatrix}
 1.96 \\
 -3.33\\
0.41
\end{pmatrix}=\begin{pmatrix}
 0.59 \\
 -1 \\
0.12
\end{pmatrix}"



5th Iteration


"D^-x_4=\begin{pmatrix}
 -2&1 & 1 \\
 5&-1& -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 0.59 \\
 -1\\
0.12
\end{pmatrix}=\begin{pmatrix}
 -2.06 \\
 3.58 \\
-0.47
\end{pmatrix}"




"x_5=\frac{1}{3.58}\begin{pmatrix}
 -2.06 \\
 3.58 \\
-0.47
\end{pmatrix}=\begin{pmatrix}
 -0.57 \\
 1 \\
-0.13
\end{pmatrix}"



6th Iteration


"D^{-1}x_5=\begin{pmatrix}
 -2&1 & 1 \\
 5&-1& -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 -0.57 \\
 1 \\
-0.13
\end{pmatrix}=\begin{pmatrix}
 2.02 \\
 -3.48 \\
0.44
\end{pmatrix}"



"x_6=\frac{1}{3.48}\begin{pmatrix}
 2.02\\
 -3.48 \\
0.44
\end{pmatrix}=\begin{pmatrix}
 0.58 \\
 -1 \\
0.13
\end{pmatrix}"



8th Iteration


"D^{-1}x_7=\begin{pmatrix}
 -2&1 & 1 \\
 5&-1 & -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 -0.58 \\
 1\\
-0.13
\end{pmatrix}=\begin{pmatrix}
 2.03 \\
 -3.5 \\
0.45
\end{pmatrix}"




"x_8=\frac{1}{3.5}\begin{pmatrix}
 2.03 \\
 -3.5 \\
0.45
\end{pmatrix}=\begin{pmatrix}
 0.58 \\
 -1 \\
0.13
\end{pmatrix}"



9th Iteration


"D^{-1}x_8=\begin{pmatrix}
 -2&1& 1\\
 5&-1 & -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 0.58 \\
 -1 \\
0.13
\end{pmatrix}=\begin{pmatrix}
 -2.03 \\
 3.51 \\
-0.45
\end{pmatrix}"



"x_9=\frac{1}{3.51}\begin{pmatrix}
 -2.03 \\
 3.51 \\
-0.45
\end{pmatrix}=\begin{pmatrix}
 0.58 \\
 1 \\
-0.13
\end{pmatrix}"



10th Iteration


D"^-x_9=\begin{pmatrix}
 -2&1 & 1\\
 5&-1 & -3\\
-1&0&1
\end{pmatrix}\begin{pmatrix}
 0.58 \\
 -1 \\
0.13
\end{pmatrix}=\begin{pmatrix}
 2.03 \\
 -3.51\\
0.45
\end{pmatrix}"



"x_{10}=\frac{1}{3.51}\begin{pmatrix}
 2.03 \\
 -3.51 \\
0.45
\end{pmatrix}=\begin{pmatrix}
 0.58 \\
 -1\\
0.13
\end{pmatrix}"




The dominant eigenvalue of D- "3.51"


The magnitude of the smallest eigenvalues of "D=\frac{1}{3.51}=0.2849"

and the corresponding eigenvector of


"D=\begin{pmatrix}
 0.58\\
 -1 \\
0.13
\end{pmatrix}"